OpenStudy (anonymous):

What is the value of the expression? cos-1 (√2)/2

OpenStudy (solomonzelman):

Hint:

OpenStudy (solomonzelman):

\(\large\color{blue}{ \cos^{-1} \frac{\sqrt{2} } {2}= \text{__}° }\)